OpenStudy (anonymous):

Trig problem. I understand how to find equivalent angles in a circle using the three common traingles. However if I have an arbitrary number ( (5pi)/9 ) how would I find an equivalent for this number? I know the answer. I just want to know HOW? thanks.

zepdrix (zepdrix):

A full rotation around the unit circle is a distance of 2pi radians. So if you started at some arbitrary point (5pi/9), You could go an entire revolution around the circle and end up at the same point. (5pi/9) + 2pi = Your new angle. (You need to get a common denominator to combine them). More generally, you could write this with ANY number of rotations, 2pi(k)= K full rotations around the circle. \[\theta \pm 2 \pi k\] Is this what you were asking about? :) Or were you asking about converting from radians to degrees? +_+

OpenStudy (anonymous):

Sorry if my questionis confusing, haha.. I mean. If 5pi/9 falls into quadrant I what would be the equivalent for quadrant II, III, IV. like this. How could I find it?

zepdrix (zepdrix):

Oh for equivalent angles in different quadrants, what you are actually doing is rotating a QUARTER of a circle :) So instead of adding 2pi, let's add 1/4 of that. \[\theta + \frac{ \pi }{ 2 }k\]

OpenStudy (anonymous):

Thank you very much for explaining that! :)